PHYS2ll4 Spring 2017 (5) Two solenoids are of a total of 2000 turms of wire. copper m made of When currents are present in the Solenoid I 2 mm two is 1 to the B. 2 axis solenoid ratio the while along is of II, of magnetic B the is field 1/4 (6) A square loop wire lies in the plane the page. A uniform page. When of of B. the square loop starts to rotate constant fiel just about the as magnetic The is into the current in loop is the C. Counterclockwise D. E None of two of the loop sides and counterclockwise in other two the above (7) The which of these possible orientations of magnetic ina uniform field B. magnetic torque the dipole magnetic (8) A of its loop Wire is A, sides at frequency f The induced emf is a maximum when: B. the flux is half its maximum value c, the flux is a maximum the flux is zero to unifom magnetic field and then spun around E. the of the flux with respect to time is zero one of the above
Answer #1

5) D) 1/2

BI = mue*N*I/L1

BII = mue*N*I/L2

BI/BII = L2/L1

= 1/2

6) B. clockwise

7) A

8) C) the flux is zero.